# mpdframe frametitlebackgroundcolor =teal!15, frametitlerule =true, roundcorner =1pt, middlelinewidth =1pt, innermargin =0.1cm, outermargin =0.1cm, innerleftmargin =0.1cm, innerrightmargin =0.1cm, innertopmargin =0.1cm, innerbottommargin =0.1cm

### v3

@@ 1 Introduction
Despite their advantages, using mocks within tests requires considerable engineering effort. While designing the tests for one unit, developers must first identify the components that may be replaced with mocks [3]. They must also determine how these mocks behave when triggered with a certain input, i.e., how they are stubbed [4, 5]. In order to address these challenges, several techniques have been developed to automatically generate mocks. For example, mocks have been generated through search-based algorithms by including contextual information, such as the state of the environment, in the search space for input data generation [6, 7]. Dynamic symbolic execution has been used to define the behavior of mocks through generated mock classes [8]. System test executions can be monitored to derive unit tests that use mocks [9]. This preliminary research has validated the concept of mock generation. However, the test inputs and the mocks produced by these techniques are either synthetic or manually written by developers per their assumptions of how the system behaves. These approaches do not guarantee that the generated mocks reflect realistic behaviors as observed in production contexts.
@@
  • An evaluation of the approach on $33$ widely-used, large, open-source Java applications, demonstrating the feasibility of generating mocks on real world software.
@@ 3.3.1 Identification of Test Generation Targets
It is not possible to generate test cases with mocks for all methods with nested method calls. For example, a static method invoked within another method is typically not mocked [3]. Also, it is not feasible to replace an object created within the body of a method, and subsequently mock the interactions made with it. Therefore, rick defines a set of rules to determine the methods that can be valid targets for the generation of test cases and mocks. Developers can provide an initial set of methods of interest for test generation, which rick can consider. Otherwise, by default, rick analyzes the whole application to determine the set of valid candidates.
@@
Second, rick identifies the nested method calls within each of the selected MUTs which could be mocked. An MUT may have several nested method calls, not all of which are suitable for mocking. For it to be mocked, a nested method call must be invoked on a parameter or a field, such that a mock can be injected to substitute it in the generated test. Next, the declaring type of this parameter or field should be different from the type of the MUT, in order to represent an interaction of the MUT with an external type, per the theory of mocking external resources. rick stubs methods that return a primitive or String value. Mocks are never returned from stubbed methods. If a nested method call fulfills all these criteria, rick identifies it as a mockable method call.
@@ 3.5 Implementation
rick is implemented in Java. MUTs and their corresponding mockable method calls are identified through static analysis with Spoon [28]. Once identified, they are instrumented and monitored in production with Glowroot, an open-source Application Production Monitoring agent [^3] https://glowroot.org/. Glowroot is a well-documented, industry-grade tool. It has a low overhead and is stable. This makes it the best fit for monitoring production executions for mock generation. Data collection from production is handled through serialization by XStream [^4] https://x-stream.github.io/. rick relies on the code generation capabilities of Spoon [^5] https://spoon.gforge.inria.fr/ to produce JUnit tests [^6] https://junit.org/. These tests define and use Mockito [^7] https://site.mockito.org/ mocks. rick uses some capabilities provided by the pankti framework [15].
@@ 4 Experimental Methodology
This section introduces our experimental methodology. We first describe the open-source projects we use as case studies to evaluate mock generation with rick. Then, we describe the production conditions that we use to collect data for test generation. Next, we present the research questions we aim to answer through this evaluation, and define the protocol that we use to do so.
@@ 4.1 Case Studies
Our first case study is the web-based routing application based on OpenStreetMap called GraphHopper  [^8] https://www.graphhopper.com/open-source/. It allows users to find the route between locations on a map, considering diverse means of transport and other routing information such as elevation. We use version 5.3 of GraphHopper, with 89898989K lines of code (LOC), 5,84458445,8445 , 844 commits, and over 4444K methods. The GraphHopper project repository on GitHub has 3.73.73.73.7K stars.
@@
The last case study is PDFBox, a PDF manipulation command-line tool developed and maintained by the Apache Software Foundation [^10] https://pdfbox.apache.org/ [31]. Some of the functionalities offered by PDFBox include the extraction of text and images from PDF documents, conversion between text files and PDF documents, encryption and decryption, and splitting and merging of PDF documents. As highlighted in Table I, we use version 2.0.24 of PDFBox, which has 165165165165K LOC, over 9999K methods, 8,79787978,7978 , 797 commits, and has been starred by 1.71.71.71.7K GitHub users.
@@ 4.2 Production Usage
TABLE II: Characteristics of the workloads for the case studies in production
@@
Once the candidate MUTs for an application are identified, the instrumented application is deployed and run under a certain workload. At runtime, rick collects data, which it will use for test and mock generation. As rick aims at consolidating test suites for mission-critical functionalities that everybody relies on, we design workloads that exercises common features and operations. We manually analyze the case studies’ codebase and refer to their documentation, in order to design workloads that capture commonly used features and operations.
Table II summarizes the key characteristics of the workloads that represent typical usage scenarios for the three applications. Rows 2 and 3 capture the scope for test generation that we consider for our experiments. The number of candidate MUTs actually invoked in production is indicated by MUT_INVOKED, while MOCKABLE_INVOKED represents the number of distinct mockable methods invoked within these invoked MUTs. We also report the number of times the MUTs and their mockable methods are invoked in the rows MUT_INVOCATIONS and MOCKABLE_INVOCATIONS, respectively. The number of invocations of MUTs and mockable methods demonstrate the relevance and comprehensiveness of the production scenarios we design. They represent the extent to which we exercise the three applications in production, and reflect actual usage of their functionalities. In total, rick focuses on 212 target methods, that invoke 199 mockable method calls. Our experiments trigger more than 7.5 million invocations of these 212 methods.
@@ 4.4 Protocols for Evaluation
Protocol for RQ5: RQ5 is a qualitative assessment of the tests generated by rick. It serves as a proxy for the readiness of the rick tests to be integrated into the test suite of projects. To assess the quality of the generated tests, we conduct a developer survey, presenting a set of $66$ tests generated by rick for $22$ MUTs in GraphHopper, to $55$ software testers from the industry. The tests that we present to the survey participants are carefully selected to represent all kinds of oracles. Developer surveys have previously been conducted to assess mocking practices [3, 12]. The key novelty of our survey consists in assessing mocks that have been automatically generated.
@@ 17 InOrder orderVerifier = inOrder(mockDataAccess);
We now discuss cases where mutants are not killed by rick tests. First, a mutant will be undetected if it results in a behavior that is equivalent to that of the original MUT. This is a well-known limitation of mutation analysis [41]. Second, a mutant will be undetected if the production input cannot infect the program’s state, or if a mock-based oracle does not capture its side-effects. The challenges of producing test cases that can effectively infect, propagate and observe the effects of a mutant is well known in the literature [42, 43]. An interesting prospect for future work is to trigger the detection of alive covered mutants through stubbing. For example, we observe this for mutant #1 in loadExisting (LABEL:lst:loadExisting-mutants), which is undetected by the OO in LABEL:lst:loadExisting-oo. The mutation causes the MUT to directly return true, which is indeed the expected value specified by the OO. However, this mutant is killed by PO and CO, as the mock method calls they specify no longer occur within the MUT. Similarly, the CO in LABEL:lst:loadExisting-co does not kill mutants #4 and #5 in LABEL:lst:loadExisting-mutants. The method calls specified by the CO still occur with the same frequency and in the expected order, but with different parameters. Parameter verification is not the focus of the CO, but the PO in LABEL:lst:loadExisting-po detects the mutants and fails.
@@
[style=mpdframe,nobreak=true,frametitle=Answer to RQ4] All three types of oracles are effective at detecting regressions introduced within MUTs, and OO, PO, and CO can detect bugs. Moreover, $1616$, $1818$, and $22$ mutants are only killed by OO, PO, and CO, respectively. The rick tests with mock-based oracles complement each other for regression testing, making them a useful addition to the test suite.
@@ 5.5 Results for RQ5 [Quality]
We begin each meeting by introducing the concepts and terminology of mock-based testing, the rick pipeline, and our experiments with GraphHopper. In order to demonstrate the features of the tests generated by rick, we select a total of $66$ generated tests, one for each of the three mock-based oracles for $22$ MUTs defined in GraphHopper. We select the first two MUTs from Table III that meet the selection criteria mentioned in subsection 4.4. The two selected MUTs, MUT#16 and MUT#6, have $5858$ and $1313$ LOC, respectively. The tests generated for MUT#16 have two stubs, and a method call on an external parameter mock object. The tests for MUT#6 have one stub and three method calls on a mocked field. Excluding comments, the number of lines of code across the six generated tests is 39 (median 6 lines of code for each test). We introduce the two MUTs, MUT#6 and MUT#16, to the participant, also presenting their source code. We invite them to clone a fork of GraphHopper which includes the generated tests. During the meeting, we browse through the generated tests with them via screen sharing. Finally, we ask the participant three sets of questions about the generated tests while documenting their responses. These questions relate to mocking effectiveness, i.e., how mocks are used within the tests, as well as the structure and understandability of the generated tests.
@@
[style=mpdframe,nobreak=true,frametitle=Answer to RQ5] Five experienced developers confirm that the concept of data collection in production is relevant for the generation of tests with mocks. They all appreciate the systematic “Arrange-Act-Assert" template for the test which contributes to the overall good understandability of the generated tests.
@@ 6.2 Threats & Limitations
Application Vs Library A source of threat to the external validity of our findings arises from the software projects we consider for the evaluation of rick. We make sure that the three projects are 1) complex, and 2) from diverse domains. Indeed, we consider a library, a desktop application, and a backend application. Yet, we do not guarantee that our findings hold for applications and libraries in other languages such as TypeScript, Rust, or nim. We do believe it is the case though.
Program Evolution Tests and programs co-evolve within software projects [47, 48, 49]. The tests generated by rick are no different, they may be impacted by changes in the application code. A refactoring in the source code, such as a modification in the order of method invocations, or a change in the parameters of an existing method, may require changes in the generated tests. However, if these modifications are not semantically relevant, this is tedious, low value work for developers. This limitation is shared by all regression test generation techniques. In this case, there is always the solution for developers to regenerate new tests when significant changes are made to the methods under tests.
@@ 7.2 Mock Generation
Several studies propose approaches to automatically generate mocks, albeit not from production executions. For example, search-based test generation can be extended to include mock objects [21], to mock calls to the file system [6] and the network [7]. Symbolic execution may also be used to generate mocks [22, 68, 8], to mock the file system [69], or a database for use in tests [70]. Honfi and Micskei [71] generate mocks to replace external interactions with calls to a parameterized sandbox. This sandbox receives inputs from the white-box test generator, Pex [23]. Moles by Halleux and Tillmann [72] also works with Pex to isolate the unit under test from interactions with dependencies by delegating them to alternative implementations. Salva and Blot [73] propose a model-based approach for mock generation. Bhagya et al. [74] use machine learning models to mock HTTP services using network traffic data. GenUTest [75] generates JUnit tests and mock aspects by capturing the interactions that occur between objects during the execution of medium-sized Java programs. StubCoder [5] by Zhu et al. uses an evolutionary algorithm to generate new stubs and repair incorrect stubs within existing JUnit tests. Abdi and Demeyer [76] leverage mocking within their proposed test transplantation technique that ports client tests into library test suites. ARTISAN by Gambi et al. [77] instruments end-to-end GUI tests of Android applications, in order to carve unit tests that mock classes of the Android framework.

### v4

# Mimicking Production Behavior with Generated Mocks

@@ 1 Introduction
Despite their advantages, using mocks within tests requires considerable engineering effort. Developers must first identify the components that may be replaced with mocks [3]. They must also determine how these mocks behave when triggered with a certain input, i.e., how they are stubbed [4, 5]. In order to address these challenges, several techniques have been developed to automatically generate mocks. For example, mocks have been generated through search-based algorithms by including contextual information, such as the state of the environment, in the search space for input data generation [6, 7]. Dynamic symbolic execution has been used to define the behavior of mocks through generated mock classes [8]. System test executions can be monitored to derive unit tests that use mocks [9]. This preliminary research has validated the concept of mock generation. However, the test inputs and the mocks produced by these techniques are either synthetic or manually written by developers per their assumptions of how the system behaves. These approaches do not guarantee that the generated mocks reflect realistic behaviors as observed in production contexts.
@@
  • An evaluation of the approach on $33$ widely-used, large, open-source Java applications, demonstrating the feasibility and benefits of generating mocks.
@@ 3.3.1 Identification of Test Generation Targets
It is not possible to generate test cases with mocks for all methods with nested method calls. For example, a static method invoked within another method is typically not mocked [3]. Also, it is not feasible to replace an object created within the body of a method, and subsequently mock the interactions made with it. Therefore, rick includes a set of rules to determine the methods that can be valid targets for the generation of test cases and mocks. It is also possible for developers to provide an initial set of methods of interest, which rick can consider.
@@
Second, rick identifies the nested method calls within each of the selected MUTs which could be mocked. An MUT may have several nested method calls, not all of which are suitable for mocking. For it to be mocked, a nested method call must be invoked on a parameter or a field, such that a mock can be injected to substitute it in the generated test. Next, the declaring type of this parameter or field should be different from the type of the MUT, in order to represent an interaction of the MUT with an external type, per the theory of mocking external resources. rick stubs methods that return a primitive or String value. Mocks are never returned from stubbed methods. Nested method calls that meet all these criteria, are marked as a mockable method calls.
@@ 3.5 Implementation
rick is implemented in Java. MUTs and their corresponding mockable method calls are identified through static analysis with Spoon [28]. Once identified, they are instrumented and monitored in production with Glowroot, an open-source Application Production Monitoring agent [^3] https://glowroot.org/. Glowroot is a well-documented, industry-grade tool. It has a low overhead and is stable. This makes it the best fit for monitoring production executions for mock generation. Data collection from production is handled through serialization by XStream [^4] https://x-stream.github.io/. rick relies on the code generation capabilities of Spoon [^5] https://spoon.gforge.inria.fr/ to produce JUnit tests [^6] https://junit.org/. These tests define and use Mockito [^7] https://site.mockito.org/ mocks, specifically the mockito-inline flavour, which allows mocking final classes. By default, rick generates three separate tests that contain the parameter oracle the call oracle, and the output oracle. If the MUT does not return a primitive or a String, rick generates only two tests with the parameter and the call oracles. rick uses some capabilities provided by the pankti framework [15].
@@ 4 Experimental Methodology
This section introduces our experimental methodology. We describe the open-source projects we use as case studies to evaluate mock generation with rick. Then, we describe the production conditions that we use to collect data for test generation. Next, we present our research questions and define the protocol that we use to answer them.
@@ 4.1 Case Studies
Our first case study is the web-based routing application based on OpenStreetMap called GraphHopper  [^8] https://www.graphhopper.com/open-source/. It allows users to find the route between locations on a map, considering diverse means of transport and other routing information such as elevation. We use version 5.3 of GraphHopper, with 89898989K lines of code (LOC), 5,84458445,8445 , 844 commits, and over 4444K methods. The project’s repository on GitHub has 3.73.73.73.7K stars.
@@
The last case study is PDFBox, a PDF manipulation command-line tool developed and maintained by the Apache Software Foundation [^10] https://pdfbox.apache.org/ [31]. It can extract text and images from PDF documents, convert between text files and PDF documents, encrypt and decrypt, and split and merge PDF documents. As highlighted in Table I, we use version 2.0.24 of PDFBox, which has 165165165165K LOC, over 9999K methods, 8,79787978,7978 , 797 commits, and has been starred by 1.71.71.71.7K GitHub users.
@@ 4.2 Production Usage
TABLE II: Characteristics of the workloads for the case studies in production: The four metrics are defined in subsection 4.2.
@@
Once the candidate MUTs for an application are identified, the instrumented application is deployed and run under a certain workload. As rick aims at consolidating test suites for mission-critical functionalities that everybody relies on, we design workloads that exercise common features. We manually analyze the case studies’ codebase and refer to their documentation, in order to design workloads that capture commonly used features and operations.
Table II summarizes the key characteristics of the workloads. Rows 2 and 3 capture the scope for test generation that we consider for our experiments. The number of candidate MUTs actually invoked in production is indicated by MUT_INVOKED, while MOCKABLE_INVOKED represents the number of distinct mockable methods invoked within these invoked MUTs. We also report the number of times the MUTs and their mockable methods are invoked in the rows MUT_INVOCATIONS and MOCKABLE_INVOCATIONS, respectively. The number of invocations of MUTs and mockable methods demonstrate the relevance and comprehensiveness of the production scenarios we design. They represent the extent to which we exercise the three applications in production, and reflect actual usage of their functionalities. In total, rick focuses on 212 target methods, that invoke 199 mockable method calls. Our experiments trigger more than 7.5 million invocations of these 212 methods.
@@ 4.4 Protocols for Evaluation
Protocol for RQ5: RQ5 is a qualitative assessment of the tests generated by rick. It serves as a proxy for the readiness of the rick tests to be integrated into the test suite of projects. To assess the quality of the generated tests, we conduct a developer survey, presenting a set of $66$ tests generated by rick for $22$ MUTs in GraphHopper, to $55$ software testers from the industry. We carefully select the tests to present to the survey participants in order to have a representation of the three kinds of oracles as well as diverse mocking contexts i.e., external field or parameter objects. Developer surveys have previously been conducted to assess mocking practices [3, 12]. The key novelty of our survey consists in assessing mocks that have been automatically generated.
@@ 17 InOrder orderVerifier = inOrder(mockDataAccess);
We now discuss cases where mutants are not killed by rick tests. First, a mutant will be undetected if it results in a behavior that is equivalent to that of the original MUT. This is a well-known limitation of mutation analysis [41]. Second, a mutant will be undetected if the production input cannot infect the program’s state, or if a mock-based oracle does not capture its side-effects. The challenges of producing test cases that can effectively infect, propagate and observe the effects of a mutant are well known in the literature [42, 43]. An interesting prospect for future work is to trigger the detection of alive covered mutants through stubbing. For example, we observe this for mutant #1 in loadExisting (LABEL:lst:loadExisting-mutants), which is undetected by the OO in LABEL:lst:loadExisting-oo. The mutation causes the MUT to directly return true, which is indeed the expected value specified by the OO. However, this mutant is killed by PO and CO, as the mock method calls they specify no longer occur within the MUT. Similarly, the CO in LABEL:lst:loadExisting-co does not kill mutants #4 and #5 in LABEL:lst:loadExisting-mutants. The method calls specified by the CO still occur with the same frequency and in the expected order, but with different parameters. Parameter verification is not the focus of the CO, but the PO in LABEL:lst:loadExisting-po detects the mutants and fails.
@@
[style=mpdframe,nobreak=true,frametitle=Answer to RQ4] All three types of oracles are effective at detecting regressions introduced within MUTs. Moreover, $1616$, $1818$, and $22$ mutants are only killed by OO, PO, and CO, respectively. The rick tests with mock-based oracles complement each other for regression testing, making them a useful addition to the test suite.
@@ 5.5 Results for RQ5 [Quality]
We begin each meeting by introducing the concepts and terminology of mock-based testing, the rick pipeline, and our experiments with GraphHopper. We demonstrate the features of the tests generated by rick by selecting a total of $66$ generated tests, one for each of the three mock-based oracles for $22$ MUTs defined in GraphHopper. We select the first two MUTs from Table III that meet the selection criteria mentioned in subsection 4.4. The two selected MUTs, MUT#16 and MUT#6, have $5858$ and $1313$ LOC, respectively. The tests generated for MUT#16 have two stubs, and a method call on an external parameter mock object. The tests for MUT#6 have one stub and three method calls on a mocked field. Excluding comments, the number of lines of code across the six generated tests is 39 (median 6 lines of code for each test). We introduce the two MUTs, MUT#6 and MUT#16, to the participant, also presenting their source code. We invite them to clone a fork of GraphHopper which includes the generated tests. During the meeting, we browse through the generated tests with them via screen sharing. Finally, we ask the participant three sets of questions about the generated tests while documenting their responses. These questions relate to mocking effectiveness, i.e., how mocks are used within the tests, as well as the structure and understandability of the generated tests.
@@
[style=mpdframe,nobreak=true,frametitle=Answer to RQ5] Five experienced developers confirm that the concept of data collection in production is relevant for the generation of tests with mocks. They all appreciate the systematic “Arrange-Act-Assert" template for the test which contributes to the overall good understandability of the tests generated by rick.
@@ 6.1 Performance
6.2 Subsequent Test Failures
As we demonstrate with RQ4, the potential future failures of a test generated by rick are meant to be indicative of regressions in the method under test, compared to the current behavior. However, there are two cases when a rick test would fail in the absence of a regression.
First, for mocked methods that perform non-deterministic actions, such as network calls, the stubs and mock-based oracles generated by rick reflect the responses observed in the field, by design. For instance, calls that result in valid responses (2XX) or 4XX errors, such as a 404 for a non-existent resource, will lead to the generation of useful tests with rick. However, rick will generate a test that may overfit a situation with a 5XX error. Such a response might not be reproduced within the generated test, causing it to fail.
1public float getWidthFromFont(int code) {
2 int gid = codeToGID(code);
3 float width = ttf.getAdvanceWidth(gid);
4 float unitsPerEM = ttf.getUnitsPerEm();
5 if (unitsPerEM != 1000) {
6 width *= 1000f / unitsPerEM;
7 }
8 return width;
9}
10@Test
11public void testGetWidthFromFont_OO() {
12 ...
13 // Assert
14 assertEquals(750.0, actual, 0.0);
15}
16@Test
17public void testGetWidthFromFont_PO() {
18 ...
19 // Assert
20 verify(mockTTF, atLeastOnce()).getAdvanceWidth(0);
21 verify(mockTTF, atLeastOnce()).getUnitsPerEm();
22}
23@Test
24public void testGetWidthFromFont_CO() {
25 ...
26 // Assert
27 InOrder ordVerifier = inOrder(mockTTF);
28 ordVerifier.verify(mockTTF, times(1)).getAdvanceWidth(anyInt());
29 ordVerifier.verify(mockTTF, times(1)).getUnitsPerEm();
30}
Listing 15: rick generates tests with the three mock-based oracles for MUT#27 of PDFBox, getWidthFromFont. Refactoring getWidthFromFont can impact the validity of these oracles.
1// Refactored getWidthFromFont
2public float getWidthFromFont(int code) {
3 int gid = codeToGID(code);
4 float unitsPerEM = ttf.getUnitsPerEm(); // reordered call
5 float width = ttf.calculateAdvanceWidth(gid); // renamed method
6 if (unitsPerEM != 1000) {
7 width *= 1000f / unitsPerEM;
8 }
9 return width;
10}
Listing 16: MUT#27 of PDFBox, getWidthFromFont, after refactoring. Relative to LABEL:lst:pre-refactoring, its mockable method calls getAdvanceWidth and getUnitsPerEm are reordered, and the former is renamed to calculateAdvanceWidth.
Second, the failure of a rick-generated test may be the result of a refactoring [46], and not a regression. For example, for a method under test m, the call oracle will fail if the order of the mock method calls within m is changed, while the semantics of m is preserved. Also, the parameter oracle may not hold if the arguments received by at least one mock method call within m deviate from their expected values. This can indicate a behavioral change within m, but also a refactoring of the stubbed method call. The parameter and call oracles are sensitive to refactoring changes. However, a failing output oracle implies that the output from m is unequal to the expected output, signifying a behavioral change in the public API.
Consider lines 1 to 9 of LABEL:lst:pre-refactoring, which present the method getWidthFromFont of PDFBox (MUT#27 in Table V). This MUT calls two mock methods on the field ttf of type TrueTypeFont, getAdvanceWidth (line 4) and getUnitsPerEm (line 5). The mock-based oracles generated by rick for getWidthFromFont are presented on line 15 (OO), lines 22-23 (PO), and lines 30-32 (CO). LABEL:lst:post-refactoring shows a refactored version of getWidthFromFont, with two semantically-preserving changes: 1) the mock method calls are reordered, i.e., getUnitsPerEm occurs first (line 4), and 2) getAdvanceWidth is renamed to calculateAdvanceWidth (line 5). These changes cause a failure of the CO and PO, but the OO still holds.
Despite their varying degree of sensitivity, all three mock-based oracles alert the developer of behavioral changes that are introduced in their code, which is also the goal of testing. The impact of code refactoring on mock-based oracles is an important direction for future work.
@@ 6.3 Threats & Limitations
Application Vs Library A source of threat to the external validity of our findings arises from the software projects we consider for the evaluation of rick. We make sure that the three projects are 1) complex, and 2) from diverse domains. Indeed, we consider a library, a desktop application, and a backend application. We do not guarantee that our findings hold for applications and libraries in other languages such as TypeScript, or nim. Still, we believe that rick would not let them down.
Program Evolution Tests and programs co-evolve within software projects [48, 49, 50]. The tests generated by rick are no different, they may be impacted by changes in the application code. A refactoring in the source code, such as a modification in the order of method invocations, or a change in the parameters of an existing method, may require changes in the generated tests. However, if these modifications are not semantically relevant, this is tedious, low value work for developers. This limitation is shared by all regression test generation techniques. In this case, developers can always regenerate new tests when significant changes are made to the methods under tests.
@@ 7.2 Mock Generation
Several studies propose approaches to automatically generate mocks, albeit not from production executions. For example, search-based test generation can be extended to include mock objects [21], to mock calls to the file system [6] and the network [7]. Symbolic execution may also be used to generate mocks [22, 69, 8], to mock the file system [70], or a database for use in tests [71]. Honfi and Micskei [72] generate mocks to replace external interactions with calls to a parameterized sandbox. This sandbox receives inputs from the white-box test generator, Pex [23]. Moles by Halleux and Tillmann [73] also works with Pex to isolate the unit under test from interactions with dependencies by delegating them to alternative implementations. Salva and Blot [74] propose a model-based approach for mock generation. Bhagya et al. [75] use machine learning models to mock HTTP services using network traffic data. GenUTest [76] generates JUnit tests and mock aspects by capturing the interactions that occur between objects during the execution of medium-sized Java programs. StubCoder [5] by Zhu et al. uses an evolutionary algorithm to generate new stubs and repair incorrect stubs within existing JUnit tests. On the other hand, ARUS [77] utilizes information from the execution of the test suite to detect and remove unnecessary stubbing. Abdi and Demeyer [78] leverage mocking within their proposed test transplantation technique that ports client tests into library test suites. ARTISAN by Gambi et al. [79] instruments end-to-end GUI tests of Android applications, in order to carve unit tests that mock classes of the Android framework.